OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (fools101):

Hi Use this Web site! http://www.mathway.com/problem.aspx?p=algebra let know if this help you!

OpenStudy (anonymous):

3(1 + x) = -2(x + -1) + 6 (1 * 3 + x * 3) = -2(x + -1) + 6 (3 + 3x) = -2(x + -1) + 6 3 + 3x = -2(-1 + x) + 6 3 + 3x = (-1 * -2 + x * -2) + 6 3 + 3x = (2 + -2x) + 6 3 + 3x = 2 + 6 + -2x 3 + 3x = 8 + -2x 3 + 3x + 2x = 8 + -2x + 2x 3 + 5x = 8 + -2x + 2x 3 + 5x = 8 + 0 3 + 5x = 8 3 + -3 + 5x = 8 + -3 0 + 5x = 8 + -3 5x = 8 + -3 5x = 5 x = 1

OpenStudy (anonymous):

Egad, I was beat to the point.

OpenStudy (anonymous):

thanks alot guys (:

OpenStudy (aravindg):

@TrashMan914 giving direct worked out answers is against openstudy COC Refer http://openstudy.com/code-of-conduct

OpenStudy (aravindg):

instead help the asker get to the solution